likeabossssssss (likeabossssssss):

The Buckley family drove 270 miles in 5 hours. The Fang family drove 364 miles in 7 hours. Which family traveled at the higher rate of speed and what was this rate? A. Buckley; 52 mi/h B. Fang; 52 mi/hr C. Buckley; 54 mi/h D. Fang; 54 mi/h b?

likeabossssssss (likeabossssssss):

@jim_thompson5910

rebeccaxhawaii (rebeccaxhawaii):

divide to see how many miles per hour 270/5=? and 364/7=?

likeabossssssss (likeabossssssss):

i did 270/5=54 and 367/7=52

rebeccaxhawaii (rebeccaxhawaii):

okay which one is higher?

likeabossssssss (likeabossssssss):

52 XD

likeabossssssss (likeabossssssss):

jk 54

rebeccaxhawaii (rebeccaxhawaii):

dont freak me out XD okay so who had 54?

likeabossssssss (likeabossssssss):

c lol sorry

rebeccaxhawaii (rebeccaxhawaii):

yes it would be c